Operations Report – May 21st

Crew 299 Operations Report 5-21-2024

SOL: 9

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission.

Rovers Used: None
General notes on rovers: None

Summary of Hab operations:
      WATER USE: 17 Gallons
      Water (static tank): 270.3 (using effective value from spreadsheet)
      Static tank pipe heater: off
      Static tank heater: off
      Toilet tank emptied: no

Summary of internet usage: Catch up on Earthly news, email responses, code development, entertainment

Summary of suits and radios: None used

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: More rock samples were made and analyzed. Radishes were watered and checked in on. Samples are also going to be further categorized.

Summary of RAM operations: Measuring tape used. Another pack of zip ties might be needed. The last pack in the RAM is running low.
Summary of observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site